40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 14 02:37:58 2014, Jenne, Update, LSC, No big locking news 
    Reply  Wed May 14 14:39:31 2014, Jenne, Update, SUS, New buttons on IFO_ALIGN screen Screenshot-Untitled_Window.png
Message ID: 9951     Entry time: Wed May 14 02:37:58 2014     Reply to this: 9953
Author: Jenne 
Type: Update 
Category: LSC 
Subject: No big locking news 

Tonight was mostly cleaning up some scripts, including the re-writing the restore and align scripts for the optics. 

The new script is in the same folder as the old one (/opt/rtcds/caltech/c1/medm/MISC/ifoalign/NewAlignSoft.py), but is not yet called from the align screen.  However, I'm using it in the carm up and down scripts, and it works nicely for the PRM.  I need to check that the offset value is okay for all the other optics (i.e. are they getting misaligned enough?), but then I'll have the new script called from the screen.  The new script, per Rana's suggestion, does not touch the bias sliders.  Rather, it puts an offset in the pitch filter banks in the coil driver output matrix-of-filter-banks.  Then the misalign routine turns the offset on, and the restore routine turns the offset off.  This way we have a nice ramp time, but don't have to do the weird calculation of number of steps to take as is done in the current script.  Also, the "save" functionality will be obsolete, since we're never touching the bias sliders except for actual alignment needs.

I'm not sure what changed, other than the HEPA being on lower, but the Xarm ALS was much better behaved tonight.  I was able to hang out around arm powers of ~1 for as long as I wanted. 

I didn't try to hand over to digital REFLDC, but I was trying a few times to engage the AO path.  With the CM board set to Plus, I hear hooting when MC IN2 is about 4dB.  With the CM board set to Minus, I didn't hear hooting, but I lost lock when I went from 13dB to 14dB. 

Also, I put the cables for the SR785 back to the "A" set of test points and excitation, so that I could take a closed loop transfer function.  However, I don't know where the latest working scripts to make a remote measurement are, so maybe we can take some loop measurements tomorrow.

The carm_cm_up script is good (for tonight) up to the prompt "Press enter to indicate that it is okay to turn on MC2 LSC FM8".  There are "read"s every step of the way, so it goes nice and slow, but it'll do everything for you except any last tweaks of the PRM alignment after the PRMI is locked.

ELOG V3.1.3-